Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

A fair die has sides labeled with 1, 2, 3, 4, 5, and 6 dots.

Problem Solving — algebra and arithmetic
Expert replies
by DivyaD » Fri Jan 18, 2019 4:11 am
A fair die has sides labeled with 1, 2, 3, 4, 5, and 6 dots. If the die is rolled 4 times, what is the probability that on at least one roll, the die will show a 6?

a) $$\frac{1}{6}$$
b) $$\frac{625}{1296}$$
c) $$\frac{649}{1296}$$
c) $$\frac{671}{1296}$$
d) $$\frac{2}{3}$$
Join the discussion
Source: — Quantitative Reasoning |

by Brent@GMATPrepNow » Fri Jan 18, 2019 5:52 am
DivyaD wrote:A fair die has sides labeled with 1, 2, 3, 4, 5, and 6 dots. If the die is rolled 4 times, what is the probability that on at least one roll, the die will show a 6?

a) $$\frac{1}{6}$$
b) $$\frac{625}{1296}$$
c) $$\frac{649}{1296}$$
c) $$\frac{671}{1296}$$
d) $$\frac{2}{3}$$
We want P(select at least one 6)
When it comes to probability questions involving "at least," it's best to try using the complement.
That is, P(Event A happening) = 1 - P(Event A not happening)
So, here we get: P(getting at least one 6) = 1 - P(not getting at least one 6)
What does it mean to not get at least one 6? It means getting zero 6's.
So, we can write: P(getting at least one 6) = 1 - P(getting zero 6's)

P(getting zero 6's)
P(getting zero 6's) = P(no 6 on 1st roll AND no 6 on 2nd roll AND no 6 on 3rd roll AND no 6 on 4th roll)
= P(no 6 on 1st roll) x P(no 6 on 2nd roll) x P(no 6 on 3rd roll) x P(no 6 on 4th roll)
= 5/6 x 5/6 x 5/6 x 5/6
= 625/1296

P(getting at least one 6) = 1 - 625/1296
= 671/1296

Answer: D

Cheers.
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by fskilnik@GMATH » Fri Jan 18, 2019 4:30 pm
DivyaD wrote:A fair die has sides labeled with 1, 2, 3, 4, 5, and 6 dots. If the die is rolled 4 times, what is the probability that on at least one roll, the die will show a 6?

$$ a) \frac{1}{6}\,\,\,\,\,\,\, b) \frac{625}{1296}\,\,\,\,\,\,\,c) \frac{649}{1296}\,\,\,\,\,\,\,d) \frac{671}{1296}\,\,\,\,\,\,\,e) \frac{2}{3}$$
$$? = 1 - P\left( {\underbrace {{\rm{no}}\,\,{\rm{6}}\,\,{\rm{in}}\,\,{\rm{all}}\,\,{\rm{4}}\,\,{\rm{times}}}_{{\rm{unfavorable}}}} \right)$$
$$\left. \matrix{
{\rm{Total}}\,\,:\,\,\,{6^4}\,\,{\rm{equiprobable}}\,\,{\rm{outcomes}}\,\,\, \hfill \cr
{\rm{unfav}}\,\,{\rm{:}}\,\,{5^4}\,\,{\rm{of}}\,\,{\rm{them}} \hfill \cr} \right\}\,\,\,\,\, \Rightarrow \,\,\,\,P\left( {{\rm{unfavorable}}} \right) = {\left( {{5 \over 6}} \right)^4}$$
$$? = {{{6^4} - {5^4}} \over {{6^4}}} = {{\left( {{6^2} + {5^2}} \right)\left( {6 + 5} \right)\left( {6 - 5} \right)} \over {{6^4}}} = {{61 \cdot 11} \over {{6^4}}} = {{610 + 61} \over {{6^4}}}\,\,\,\,\, \Rightarrow \,\,\,\,\,\left( {\rm{D}} \right)$$


We follow the notations and rationale taught in the GMATH method.

Regards,
Fabio.
Fabio Skilnik :: GMATH method creator ( Math for the GMAT)
English-speakers :: https://www.gmath.net
Portuguese-speakers :: https://www.gmath.com.br
Join the discussion